expansion of a square garden. each side is increased by 7 ft new total area will be 144 square ft. what is the original length of the garden.
Found 2 solutions by ash89, checkley71:
Answer by ash89(1)
Answer by checkley71(8403)

LET X BE THE ORIGINAL LENGTH OF THE SIDES OF THE SQUARE.
THEN THE NEW LENGTH IS X+7+7=X+14
(X+14)^2=144
X^2+28X+196=144
X^2+28X+196-144=0
X^2+28+52=0
(X+26)(X+2)=0
X+26=0
X=-26 NOT AN ANSWER.
X+2=X-2 NOT AN ANSWER.
FOR A SUARE AREA=144 FT^2 THE SIDES ARE 12 FT.
HOWEVER YOU ADDED 14 FT. TO THE ORIGINAL SIDE WHICH DOESN'T FIT THE ANSWER OF 144 FT^2.